# No exponential quantum speedup for $\mathrm{SIS}^{\infty}$ anymore

### v2

@@ 4.2 From zero-sum to reducible vector
For each $j\in S_{1}$, we define $\beta_{j}\in\{\pm 1\}$ such that $\alpha_{j}\beta_{j}\in B_{1}$, i.e., $\beta_{j}=1$ if $\alpha_{j}\in B_{1}$; and $\beta_{j}=-1$ if $\alpha_{1}\in B_{1}$. Now we can rewrite the $A$-nontrivial $B$-zero-sum as
@@
$L_{1}=\sum_{i\in[k]\setminus\{0,i^{*}\}}\sum_{\pi}\operatorname{sgn}(\pi)\sum_{\begin{subarray}{c}z_{1}\in S^{z_{0}}_{\pi_{1}},\ldots,z_{j-1}\in S_{\pi_{j-1}}^{z_{j-2}}\\ j:=\pi^{-1}(i^{*})\\ z_{j}\in S_{i}^{z_{j-1}}\\ z_{j+1}\in S_{\pi_{j+1}}^{z_{j}},\ldots,z_{k}\in S_{\pi_{k}}^{z_{k-1}}\end{subarray}}\alpha_{z_{j}}\beta_{z_{j}}\left(\prod_{\ell\in[k]}\beta_{z_{\ell}}\right)v_{z_{k}}.$ (via $i\notin\{0,i^{*}\}$)
@@ 4.6 Improved $\mathrm{CIS}$ algorithms
Recall that in the $\mathrm{CIS}$ problem, we are given an arbitrary set of allowed coefficients. To handle such generality, in Lemma 4.5 we will set $B_{1},\ldots,B_{k}$ as singleton sets to ensure $B_{i}-B_{i}=\{0\}$ for all $i\in[k]$. This is formalized in the following Theorem 4.14.
@@
Observe that in Theorem 4.15 we pay a cost of roughly $n^{k+1}$ to remove $2k$ elements, though these elements need to be paired up. This already hints an improvement over Theorem 1.11, as they pay a cost of roughly $n^{2k+1}$ to remove only $2k$ elements. In general, whenever we have two paired elements that need to be discarded, we only incur a linear cost, in contrast to the quadratic cost in [CLZ22].
@@
The following fact completes the edge case in Item 2 of Fact 4.16. Its proof is also presented in Appendix A. A similar result over the integers is due to Erdős and Turán [ET36].
@@
The above results fall into the richer literature of quantitative Szemerédi’s theorem. Using more advanced techniques, the bound in Fact 4.17 can be strengthened to $c\leq(1-o(1))\cdot q$ and $3$-AP can be extended into longer APs. Numerous works are along this line and we refer interested readers to [KM23, LSS24] for recent breakthroughs.
@@
Let $\overline{A}\subseteq\mathbb{F}_{q}$, which has size $1\leq c\leq q-2$. Then we have the following cases.
@@ 4.7 Further optimizations
The starting point is the $(\pm\lfloor q/2\rfloor)$-zero-sum algorithm $\mathcal{A}_{0}$ from Lemma 4.3 that uses roughly $m_{0}(n)=n$ input vectors. Then we apply Lemma 4.9 to obtain a $(\pm\lfloor q/(2k_{1})\rfloor)$-zero-sum algorithm that uses roughly $m_{1}(n)=p_{1}^{p_{1}^{2}/2}n^{p_{1}}$ input vectors. Then we apply Lemma 4.9 to divide $q_{2}$ which ends up needing roughly
@@ 5.1 On our results
Such problems naturally benefit from fast matrix multiplication algorithms. Let $2\leq\omega<2.372$ denote a constant such that $n\times n$ matrix multiplication (and inversion) over $\mathbb{F}_{q}$ can be done in deterministic classical time $n^{\omega}\operatorname{polylog}(n,q)$. See [ADW+25] and references within for latest results. The following Fact 5.1 provides what we need, and its proof can be found in Appendix B.
@@
On the other hand, tweaking our algorithms may not be necessary. There is a simple reduction, which is essentially [CLZ22, Appendix A] and is credited to Regev, that converts a $(\pm h_{1})$-zero-sum algorithm $\mathcal{A}_{1}$ over $\mathbb{Z}/q_{1}\mathbb{Z}$ and a $(\pm h_{2})$-zero-sum algorithm $\mathcal{A}_{2}$ over $\mathbb{Z}/q_{2}\mathbb{Z}$ into a $(\pm h_{1}h_{2})$-zero-sum algorithm over $\mathbb{Z}/(q_{1}q_{2}\mathbb{Z})$. The observation is that $v\equiv 0\pmod{q_{1}q_{2}}$ is equivalent to (1) $v\equiv 0\pmod{q_{1}}$ and (2) $(v/q_{1})\equiv 0\pmod{q_{2}}$. Therefore we can first run $\mathcal{A}_{1}$ to obtain zero-sums modulo $q_{1}$ (i.e., satisfying condition (1)); then run $\mathcal{A}_{2}$ on those zero-sums divided by $q_{1}$ to obtain zero-sums modulo $q_{2}$ (i.e., satisfying condition (2)). The final sample complexity is the sample complexities multiplied.
@@ 5.2 Works by Imran, Ivanyos, Sanselme, and Santha
The most comparable papers to the present one are a line of work by Ivanyos and coauthors [ISS07, IS17, II24]. These papers are all motivated by the Chevalley–Warning theorem and by designing quantum algorithms for versions of the Hidden Subgroup Problem; yet, all contain intermediate results that can be viewed as classical algorithms for $\mathrm{SIS}^{\infty}$ or $\mathrm{CIS}$.
@@
The followup work of Ivanyos and Santha [IS17] generalized the problem and algorithm to solving systems of diagonal $d$th-power equations; this is equivalent to $\mathrm{CIS}$ with $A=\{a^{d}:a\in\mathbb{F}_{q}\}$. They gave a $\operatorname{poly}(m,\log q)$-time classical algorithm whenever
@@ 5.3 Quantum algorithms from Regev’s reduction
Aside from the work of Chen, Liu, and Zhandry [CLZ22], at least two other candidate exponential quantum speedups have been derived via Regev’s framework of converting an AA-$\mathrm{CIS}$-type problem (usually in its equivalent dual version) to a quantum decoding [CT24] problem. These are: the Yamakawa–Zhandry paper [YZ24], which gives a provable exponential quantum speedup in the random oracle query model; and, the DQI paper of Jordan et al. [JSW+24], which gives a seeming exponential quantum speedup for the “OPI” problem of fitting a low-degree univariate polynomial over $\mathbb{F}_{q}$ to a range of points (see also the followup [CT24]).
@@ 5.4 $\mathrm{SIS}^{\infty}$-related problems
This form of the problem was called LIN-SAT in [JSW+24]. This viewpoint is adopted in the Yamakawa–Zhandry problem [YZ24], where $q$ is exponentially large in $m$, $\mathcal{C}$ is an efficiently decodeable, list-recoverable code (specifically, a folded Reed–Solomon code), and $A\subseteq\mathbb{F}_{q}$ is a random set of cardinality roughly $q/2$. It is also the viewpoint in the Optimal Polynomial Interpolation (OPI) problem from [JSW+24], where we work in the targeted version, $q\approx m$, $k\approx m/10$, and $\mathcal{C}$ is the Reed–Solomon code.
@@
We remark that the $q=3$ case of $\{\pm 1\}$-$\mathrm{CIS}$ is particularly attractive. In its original, primal version, it is asking to partition a sequence of $m$ vectors from $\mathbb{F}_{q}^{n}$ into two subsequences of equal sum. In its dual version, it is asking to find a codeword in a given $\mathbb{F}_{3}$-linear code with all symbols in $\{\pm 1\}$, i.e., at maximal Hamming distance $m$ from the all-zero codeword. This “ternary syndrome decoding with maximal weight” problem was carefully studied in the context of the Wave cryptosystem [DST19]. To write the problem “LWE”-style, it is equivalent to solving a system of disequations over $\mathbb{F}_{3}$,
@@
This problem of learning from disequations, which also makes sense over fields larger than $\mathbb{F}_{q}$, has been studied in several prior works, dating back to Friedl et al. [FIM+02], and developed by Ivanyos [Iva07], Arora–Ge [AG11], and Ivanyos–Prakash–Santha [IPS18].
@@ 5.6 Post-quantum cryptosystem
The [DKL+18, LDK+] CRYSTALS-Dilithium signature scheme was selected for standardization by NIST’s post-quantum cryptography initiative. One of the key hardness assumptions underlying its security is “Module-$\mathrm{SIS}^{\infty}$”, which is a variant of $\mathrm{SIS}^{\infty}$. A concrete, comparable version of $\mathrm{SIS}^{\infty}$ that is presumed intractable has $q\approx 2^{23}$, $s\sim q/8$, and $m\sim 1.9n$, with $n=1280$. Our $\mathrm{SIS}^{\infty}$ algorithms are successful at handling $q\gg n$, but require $m\geq Cn^{4}$ in order to achieve $s\sim q/8$. Again, it would be interesting to investigate subexponential-time tradeoffs that interpolate between the polynomial-time-solvable case of $m\approx n^{4}$ and the presumed exponential-time case of $m\approx n$.
@@ Acknowledgments
We are very grateful to David Gosset for his contributions to the early stages of this work. RO thanks Siddhartha Jain for helpful discussions. KW thanks Ce Jin, Qipeng Liu, and Hongxun Wu for relevant references. We acknowledge the use of Gemini and ChatGPT to search the literature and suggest proof strategies.
@@ Appendix A Ad hoc arithmetic combinatorics
We prove the arithmetic combinatorial results here. While significantly better bounds are known [KM23, LSS24] using more sophisticated techniques, we choose to provide what we need with minimal efforts as this is not the focus of our work.

### v3

@@ 4.2 From zero-sum to reducible vector
For each $j\in S_{1}$, we define $\beta_{j}\in\{\pm 1\}$ such that $\alpha_{j}\beta_{j}\in B_{1}$, i.e., $\beta_{j}=1$ if $\alpha_{j}\in B_{1}$; and $\beta_{j}=-1$ if $\alpha_{j}\in B_{1}$. Now we can rewrite the $A$-nontrivial $B$-zero-sum as
@@
The intuition of $u$ is as follows. Each layer in the tree enforces a decomposition of an $H$-zero-sum into components aligned with $H_{0},H_{1},\ldots,H_{k}$; the permutation sum with alternating signs implements a determinant-like cancellation across branches so that for any $c\in H_{i}$, the combination $c\cdot u$ decomposes into $H^{\prime}$-conforming pieces (e.g., $\pm H_{0}$ and $B_{i}-B_{i}$) while preserving $A$-nontriviality.
@@
$L_{1}=\sum_{i\in[k]\setminus\{i^{*}\}}\sum_{\pi}\operatorname{sgn}(\pi)\sum_{\begin{subarray}{c}z_{1}\in S^{z_{0}}_{\pi_{1}},\ldots,z_{j-1}\in S_{\pi_{j-1}}^{z_{j-2}}\\ j:=\pi^{-1}(i^{*})\\ z_{j}\in S_{i}^{z_{j-1}}\\ z_{j+1}\in S_{\pi_{j+1}}^{z_{j}},\ldots,z_{k}\in S_{\pi_{k}}^{z_{k-1}}\end{subarray}}\alpha_{z_{j}}\beta_{z_{j}}\left(\prod_{\ell\in[k]}\beta_{z_{\ell}}\right)v_{z_{k}}.$ (via $i\notin\{0,i^{*}\}$)
@@ 4.6 Improved $\mathrm{CIS}$ algorithms
Recall that in the $\mathrm{CIS}$ problem, we are given an arbitrary set $A$ of allowed coefficients. For convenience, one may think of $A$ as fixed in the problem description. If $A$ is also taken as an input of the problem (say, by explicit list), our reductions in this subsection still work and will have a mild $\operatorname{poly}(q)$ time overhead to find appropriate APs in $A$.
To handle such generality of $A$, in Lemma 4.5 we will set $B_{1},\ldots,B_{k}$ as singleton sets to ensure $B_{i}-B_{i}=\{0\}$ for all $i\in[k]$. This is formalized in the following Theorem 4.14.
@@
Observe that in Theorem 4.15 we pay a cost of roughly $n^{k+1}$ to remove $2k$ elements, though these elements need to be paired up. This already hints an improvement over Theorem 1.11, as they pay a cost of roughly $n^{2k+1}$ to remove only $2k$ elements. In general, whenever we have two paired elements that need to be discarded, we only incur a linear cost, in contrast to the quadratic cost in [9].
@@
The following fact completes the edge case in Item 2 of Fact 4.16. Its proof is also presented in Appendix A. A similar result over the integers is due to Erdős and Turán [13].
@@
The above results fall into the richer literature of quantitative Szemerédi’s theorem. Using more advanced techniques, the bound in Fact 4.17 can be strengthened to $c\leq(1-o(1))\cdot q$ and $3$-AP can be extended into longer APs. Numerous works are along this line and we refer interested readers to [23, 24] for recent breakthroughs.
@@
Let $\overline{B}\subseteq\mathbb{F}_{q}$, which has size $1\leq c\leq q-2$. Then we have the following cases.
@@ 4.7 Further optimizations
The starting point is the $(\pm\lfloor q/2\rfloor)$-zero-sum algorithm $\mathcal{A}_{0}$ from Lemma 4.3 that uses roughly $m_{0}(n)=n$ input vectors. Then we apply Lemma 4.9 to obtain a $(\pm\lfloor q/(2k_{1})\rfloor)$-zero-sum algorithm that uses roughly $m_{1}(n)=q_{1}^{q_{1}^{2}/2}n^{q_{1}}$ input vectors. Then we apply Lemma 4.9 to divide $q_{2}$ which ends up needing roughly
@@ 5.1 On our results
Such problems naturally benefit from fast matrix multiplication algorithms. Let $2\leq\omega<2.372$ denote a constant such that $n\times n$ matrix multiplication (and inversion) over $\mathbb{F}_{q}$ can be done in deterministic classical time $n^{\omega}\operatorname{polylog}(n,q)$. See [2] and references within for latest results. The following Fact 5.1 provides what we need, and its proof can be found in Appendix B.
@@
On the other hand, tweaking our algorithms may not be necessary. There is a simple reduction, which is essentially [9, Appendix A] and is credited to Regev, that converts a $(\pm h_{1})$-zero-sum algorithm $\mathcal{A}_{1}$ over $\mathbb{Z}/q_{1}\mathbb{Z}$ and a $(\pm h_{2})$-zero-sum algorithm $\mathcal{A}_{2}$ over $\mathbb{Z}/q_{2}\mathbb{Z}$ into a $(\pm h_{1}h_{2})$-zero-sum algorithm over $\mathbb{Z}/(q_{1}q_{2}\mathbb{Z})$. The observation is that $v\equiv 0\pmod{q_{1}q_{2}}$ is equivalent to (1) $v\equiv 0\pmod{q_{1}}$ and (2) $(v/q_{1})\equiv 0\pmod{q_{2}}$. Therefore we can first run $\mathcal{A}_{1}$ to obtain zero-sums modulo $q_{1}$ (i.e., satisfying condition (1)); then run $\mathcal{A}_{2}$ on those zero-sums divided by $q_{1}$ to obtain zero-sums modulo $q_{2}$ (i.e., satisfying condition (2)). The final sample complexity is the sample complexities multiplied.
@@ 5.2 Works by Imran, Ivanyos, Sanselme, and Santha
The most comparable papers to the present one are a line of work by Ivanyos and coauthors [19, 20, 17]. These papers are all motivated by the Chevalley–Warning theorem and by designing quantum algorithms for versions of the Hidden Subgroup Problem; yet, all contain intermediate results that can be viewed as classical algorithms for $\mathrm{SIS}^{\infty}$ or $\mathrm{CIS}$.
@@
The followup work of Ivanyos and Santha [20] generalized the problem and algorithm to solving systems of diagonal $d$th-power equations; this is equivalent to $\mathrm{CIS}$ with $A=\{a^{d}:a\in\mathbb{F}_{q}\}$. They gave a $\operatorname{poly}(m,\log q)$-time classical algorithm whenever
@@ 5.3 Quantum algorithms from Regev’s reduction
Aside from the work of Chen, Liu, and Zhandry [9], at least two other candidate exponential quantum speedups have been derived via Regev’s framework of converting an AA-$\mathrm{CIS}$-type problem (usually in its equivalent dual version) to a quantum decoding [8] problem. These are: the Yamakawa–Zhandry paper [35], which gives a provable exponential quantum speedup in the random oracle query model; and, the DQI paper of Jordan et al. [22], which gives a seeming exponential quantum speedup for the “OPI” problem of fitting a low-degree univariate polynomial over $\mathbb{F}_{q}$ to a range of points (see also the followup [8]).
@@ 5.4 $\mathrm{SIS}^{\infty}$-related problems
This form of the problem was called LIN-SAT in [22]. This viewpoint is adopted in the Yamakawa–Zhandry problem [35], where $q$ is exponentially large in $m$, $\mathcal{C}$ is an efficiently decodeable, list-recoverable code (specifically, a folded Reed–Solomon code), and $A\subseteq\mathbb{F}_{q}$ is a random set of cardinality roughly $q/2$. It is also the viewpoint in the Optimal Polynomial Interpolation (OPI) problem from [22], where we work in the targeted version, $q\approx m$, $k\approx m/10$, and $\mathcal{C}$ is the Reed–Solomon code.
@@
We remark that the $q=3$ case of $\{\pm 1\}$-$\mathrm{CIS}$ is particularly attractive. In its original, primal version, it is asking to partition a sequence of $m$ vectors from $\mathbb{F}_{q}^{n}$ into two subsequences of equal sum. In its dual version, it is asking to find a codeword in a given $\mathbb{F}_{3}$-linear code with all symbols in $\{\pm 1\}$, i.e., at maximal Hamming distance $m$ from the all-zero codeword. This “ternary syndrome decoding with maximal weight” problem was carefully studied in the context of the Wave cryptosystem [10]. To write the problem “LWE”-style, it is equivalent to solving a system of disequations over $\mathbb{F}_{3}$,
@@
This problem of learning from disequations, which also makes sense over fields larger than $\mathbb{F}_{q}$, has been studied in several prior works, dating back to Friedl et al. [14], and developed by Ivanyos [21], Arora–Ge [3], and Ivanyos–Prakash–Santha [18].
@@ 5.6 Post-quantum cryptosystem
The [12, 26] CRYSTALS-Dilithium signature scheme was selected for standardization by NIST’s post-quantum cryptography initiative. One of the key hardness assumptions underlying its security is “Module-$\mathrm{SIS}^{\infty}$”, which is a variant of $\mathrm{SIS}^{\infty}$. A concrete, comparable version of $\mathrm{SIS}^{\infty}$ that is presumed intractable has $q\approx 2^{23}$, $s\sim q/8$, and $m\sim 1.9n$, with $n=1280$. Our $\mathrm{SIS}^{\infty}$ algorithms are successful at handling $q\gg n$, but require $m\geq Cn^{4}$ in order to achieve $s\sim q/8$. Again, it would be interesting to investigate subexponential-time tradeoffs that interpolate between the polynomial-time-solvable case of $m\approx n^{4}$ and the presumed exponential-time case of $m\approx n$.
@@ Acknowledgments
We are very grateful to David Gosset for his contributions to the early stages of this work. RO thanks Siddhartha Jain for helpful discussions. KW thanks Ce Jin, Qipeng Liu, and Hongxun Wu for relevant references. We acknowledge the use of Gemini and ChatGPT to search the literature and suggest proof strategies. We thank anonymous reviewers for helpful comments.
@@ Appendix A Ad hoc arithmetic combinatorics
We prove the arithmetic combinatorial results here. While significantly better bounds are known [23, 24] using more sophisticated techniques, we choose to provide what we need with minimal efforts as this is not the focus of our work.